OpenStudy (anonymous):

Which mathematical sentence represents the solution for a in the equation 7b = bc + 3a?

OpenStudy (anonymous):

7b-bc =3a. a = b(7-c)/3

OpenStudy (anonymous):

7b -bc = 3 a = b 7 - c/3